# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 2.20 Multiscale and multifidelity models
Here, data resolution is essential to data-driven models, impacting their accuracy, interpretability, and application. Low-resolution data is often more aggregated or sparse and may not capture slight changes in the underlying phenomenon. In this situation, models trained on low-resolution data may oversimplify complicated connections, resulting in restricted prediction ability and missing critical details. High-resolution data, on the other hand, offers extensive information, catching nuanced patterns and variations within the dataset, which can produce more accurate forecasts and identify subtle linkages. They may, however, be more computationally intensive and prone to overfitting if not appropriately managed. Despite limitations on low-resolution data, it can still be useful, particularly in situations where high-resolution data is impracticable or prohibitively expensive to gather. It is frequently utilized in large-scale environmental monitoring, satellite image processing, and some healthcare applications.
@@ 3.1 Ecology and epidemiology
The classical Fisher-KPP model and its extensions have been used in many applications, including cell biology, high energy physics, statistical physics, chemistry, ecology, and epidemiology el2019 . Along with the Fisher-KPP model, different types of nonlocal models have been discussed in this article, and we have summarised their applications in this subsection.
@@
In ecological systems, individuals within a population often interact nonlocally due to resource competition, predator-prey interactions, or spatial dispersal. Nonlocal models have been applied to study population dynamics, species coexistence, and spatial patterns in ecological communities. The classical Fisher-KPP equation and its extensions describe the space-time evolution of population densities skellam1951 ; levin2003 ; shigesada1995 ; steele1998 . The nonlocal Fisher-KPP equation (3) is called the “competitive Lotka-Volterra model” meszena2005 ; pigolotti2007 ; perthame2008 . It is a simple nonlocal model with many applications, e.g., nonlocal intra-specific competition in the predator-prey model pal2018bmb ; banerjee2017ec ; pal2019mbe ; pal2019ijbc ; pal2020mbe . As mentioned earlier, the homogeneous solution $u=1$ can be unstable due to nonlocal interaction. For this, we linearize the nonlocal model (3) around the homogeneous solution $u=1$, and then applying the Fourier transform on both sides, we obtain the eigenvalue equation
${}\lambda=-dk^{2}-\widehat{\phi}(k),$ (37)
@@
where $k$ is the wave number and $\widehat{\phi}(k)$ is the Fourier transform of the kernel function $\phi$ volpert2009 ; dornelas2021 . If we choose the kernel function as the Dirac $\delta$-function, then the nonlocal model (3) reduces to the local model (2), and the homogeneous solution $u=1$ becomes stable as $\lambda<0$ for all positive $k$ and $d$. The eigenvalue $\lambda$ may not maintain the negative sign for the other choices of the kernel functions. For example, suppose $\phi$ is a top-hat kernel function pal2018bmb ; genieys2006 , then the eigenvalue equation (37) becomes
${}\lambda=-dk^{2}-\frac{\sin(k\delta)}{k\delta},$
@@
and it is not always negative for all $k>0$; rather $\lambda$ can make positive for some $k>0$ by choosing some specific value of $\delta$ and $d$ genieys2006 . For such choices of $\delta$ and $d$, the homogeneous solution $u=1$ becomes unstable, and the nonlocal model produces a non-homogeneous stationary solution. This type of linear stability analysis around the non-trivial homogeneous solution(s) predicts the existence of heterogeneous solution(s) for the nonlocal model, and this method is known as Turing instability analysis amturing . This technique can be applied to the other nonlocal models described in this review article if the homogeneous solution satisfies the Turing instability conditions jdmurray .
@@
The considered nonlocal model with top-hat kernel function has important characteristics. For a fixed value of $\delta$, we may find the critical values $k_{c}$ and $d_{c}$ by which $\lambda=0$ holds for unique $k=k_{c}$ and $d=d_{c}$, and $\lambda>0$ holds for a range of values of $k$ with $d<d_{c}$. However, these critical values do not exist for small values of $\delta$ as $\lambda$ negative for $\delta$ approaches 00. Hence, $\delta$ (range of nonlocal interaction) requires a minimum value to unstable the homogeneous solution $u=1$ pal2018bmb ; pal2019mbe . In addition, two length scales can be identified in the presence of the top-hat kernel function: (i) the diffusion parameter $d$ as activation and (ii) the range of nonlocal interaction $\delta$ as inhibition. We can observe a Turing instability for $d\ll\delta$ (short-range activation and long-range inhibition), and it produces a Turing pattern amturing ; meinhardt1982 ; fuentes2004 . On the other hand, for $\delta\ll d$ (long-range activation and short-range inhibition), travelling wavefront propagation can be observed. In particular, for $\delta\rightarrow 0$, the nonlocal model reduces to the local classical Fisher-KPP equation and exhibits a travelling wave solution jdmurray ; berestycki2007 . The same type of behaviours can be observed for other truncated kernel functions (triangular, parabolic, etc.), but the minimum range of the nonlocal interaction varies for different functions pal2019mbe .
@@
In plant communities, the local competition term could represent the competition for light, assuming the canopy size is negligible compared to the range of the root system maciel2021 . If the opposite holds, nonlocal competition comes into the picture granato2019 ; schenk2006 . In this case, individuals compete to access the resources, which can redistribute itself furter1989 ; gourley2000 . In ecology, nonlocal models can capture many factors; e.g., Cantrell and Cosner cantrell1996 constructed a model for ladybirds feeding on aphids. They focused on the nonlocal RD model for aphid density in a particular patch that involves the whole population.
@@
Another type of solution called ‘travelling wave solution’ is studied for a partial differential equation that propagates with a constant speed while maintaining its shape in space. These solutions represent wave-like patterns that propagate through space while maintaining their shape and speed. Travelling wave solutions are common in biology, and numerous models in population dynamics produce biological waves. In the context of population dynamics, the travelling wave manifests itself as a wave of change in population density through a habitat, e.g., a plague that travels through a continent. This travelling wave solution also signifies the invasion of the species over the habitat. The travelling wave solutions have been studied for the nonlocal models, e.g., let us consider an extension of the nonlocal FKPP model (3) as billingham2003 :
@@
The initial condition plays a crucial role in finding the travelling wave solutions. A small population is introduced in a localized space and is considered the initial condition for finding the travelling wave solutions. In addition, the boundary conditions for this nonlocal model are considered as $u_{x}(0,t)=0$ and $u\rightarrow 0$ for $x\rightarrow\infty$ billingham2003 . For the “thin-tailed” kernel function $\psi(z)=\exp{(-|z|)}/2$, the nonlocal model (38) can be converted to the coupled equations billingham2003 :
@@
with localized initial conditions for the species $u$ and boundary conditions $u_{x}(0,t)=0$, $u\rightarrow 0$, $w_{x}(0,t)=0$, and $w\rightarrow 0$ for $x\rightarrow\infty$ billingham2003 . Sometimes, this conversion technique can be applied to the nonlocal model with space-time nonlocality pal2019mbe , e.g., the single species model considered in (5). Along with travelling wave solutions, periodic travelling waves (or wavetrain) are also observed in local and nonlocal models, which are the periodic solutions of travelling waves. These travelling wave solutions also signify the migration of two or more sub-populations.
The single species nonlocal model (4) in population dynamics was introduced by Britton britton1989 ; britton1990 , and it is an advanced version of the local FKPP model as it can be obtained by considering $\phi$ as the delta function. The nonlocal model (4) includes local aggregation that takes care of animal grouping for protective measures against predation, which could arise for many reasons, such as grassland herds, schools of fish or flocks of birds. Sometimes, species take time to move from one position to another. In this case, the nonlocal model (5) containing a spatio-temporal average weighted toward the current time and position better fits these types of scenarios britton1990 . This nonlocal model accounts for the influence of past conditions on present ecological processes.
Nonlocal interactions are important in multi-body systems because each body’s behaviour is influenced by its nearby neighbours and bodies at a distance. Such interactions are common in a wide range of physical and biological systems, including granular materials, colloidal suspensions, and biomolecular systems. In the case of population biology, species sense their surroundings to make different decisions, called nonlocal sensing, and these happen on different scales eftimie2007 ; levine2000 ; mogilner1999 . This nonlocal sensing affects the spatial distributions of populations and leads to species aggregation, segregation, or complex patterns bastille2018 ; benhamou2014 ; martinez2020 ; potts2014 . Researchers have studied these kinds of biological problems mathematically by aggregation-diffusion equations (34) and (35) for single- and multi-species models, respectively. In addition, the nonlocal multi-species advection-diffusion model (35) can capture a wide variety of patterns giunta2021 ; potts2019 . These types of models have been applied to prey-predator interactions fagioli2019 , cell-sorting painter2015 , and animal territoriality potts2016 . Authors in giunta2021 showed that the inclusion of self-attractive terms causes oscillatory solutions for the two-species model, and a qualitative effect on patterns occurs due to a change in the width of the spatial average.
@@
The nonlocal dispersal used in the model (8) is important for plants and animals. It characterizes the movements of organisms in a habitat. For plants, this nonlocal dispersal happens through the seed dispersal process. In the review article, Bullock et al. fitted 11 types of probability density functions as dispersal kernel functions for 144 plant species, and all of them are good fits to the grouped data sets bullock2017 . In addition, the dispersal-competition models have been used in the integrodifference equations, which are discrete in time and continuous in space allen1996 . For instance, the integrodifference equations of a population size $n(t,x)$ at time $t$ and position $x$ can have the form allen1996 :
@@
where $\phi(x,y)$ is the kernel function that satisfies the normality condition over the spatial domain $R$. In the absence of dispersal, the growth equation is of the form $n(t+1,x)=f(n(t,x))$. These models apply to the development and spread of seed-dispersed plants. Moreover, a similar model can be considered for the single-species annual plant model when seeds do not survive for more than one year allen1996 . In addition, the nonlocal dispersal is also applied to study the metapopulation dynamics. In hanski1998 , the author studied the effects of nonlocal dispersal on population persistence and the conservation of species in fragmented habitats where local populations are interconnected through dispersal, colonization, and extinction processes. It has implications for ecosystem resilience, invasive species management, and conservation biology. As research in ecology and population dynamics progresses, the study of nonlocal dispersal will continue to shed light on the intricate processes that shape biodiversity and ecological patterns.
@@
In hastings2005 , Hastings et al. provide insights into the spatial spread of invasions, a specific example of nonlocal dispersal in ecology. The paper discusses theoretical developments and empirical evidence related to the spatial dynamics of invasive species, highlighting the significance of nonlocal dispersal in driving the spread of invasive populations. Furthermore, in okubo2001 , the authors delve into the role of diffusion in ecological problems. It covers various aspects of dispersal and diffusion in ecology, including nonlocal dispersal and its effects on population dynamics and spatial ecology. The book provides a comprehensive and interdisciplinary perspective on diffusion in ecological systems and is an excellent resource for those interested in applying nonlocal dispersal in biology.
@@
These applications demonstrate the versatility of fractional nonlocal models in capturing the complexities of biological systems. Fractional calculus provides a powerful framework to describe anomalous transport, long-range interactions, and memory effects, offering valuable insights into various biological processes. For instance, fractional-order ordinary differential equations (FODEs) are used to study the memory effect on population growth or some inherent randomness at a microscopic level that regulates the macroscopic dynamics yin2018 . In ecology, FODEs have been studied in prey-predator models to account for the interactions of the population densities dependent on a certain time in the past javidi2013 ; rihan2015 ; matouk2016 . The nonlocal FODE model (10) describes the emergence and evolution of biological species in the presence of memory. Furthermore, the fractional integral representation is used in the statistical thermodynamics of confined systems to obtain generalized transport equations, distribution functions, and correlation functions that take into account the particular characteristics of confinement-induced dynamics sisman2021 .

### v2

@@ 2.19 Multiscale and multifidelity models
Data resolution is crucial for data-driven models as it affects their accuracy, interpretability, and application. Low-resolution data is often more aggregated or sparse and may not capture slight changes in the underlying phenomenon. In this situation, models trained on low-resolution data may oversimplify complicated connections, resulting in restricted prediction ability and missing critical details. High-resolution data, on the other hand, offers extensive information, catching nuanced patterns and variations within the dataset, which can produce more accurate forecasts and identify subtle linkages. They may, however, be more computationally intensive and prone to overfitting if not appropriately managed. Despite limitations on low-resolution data, it can still be useful, especially where high-resolution data is impracticable or prohibitively expensive to gather. It is frequently utilized in large-scale environmental monitoring, satellite image processing, and healthcare applications.
So far, we have discussed the models using PDEs or other continuous mathematical frameworks that model the spatial distribution of variables, such as concentrations of substances, populations, or other physical quantities. Here, we discuss the network models where systems are composed of discrete entities (nodes) and the relationships (edges) between them. Networks do not require spatial continuity and can model interactions between distant entities. Nonlocal interactions on networks in biology refer to interactions between nodes in a network (such as genes, proteins, or cells) that extend beyond their direct neighbours or immediate connections. These interactions are essential for understanding the processes and characteristics of complex biological systems in a network. This nonlocality is essential in systems where connections do not follow spatially continuous patterns, such as social networks, signalling networks, ecological food webs, and neuronal networks. While space-continuous systems and network models appear fundamentally different, they can provide complementary perspectives on complex systems. Space-continuous models are adopted in capturing continuous spatial dynamics, while network models are well-suited for nonlocal, discrete interactions.
Neurons in the brain communicate and interact through synapses, which can involve long-range connections [242, 321]. Nonlocal models have been employed to describe neurons’ collective activity and synchronization in neural networks. In addition, Alzheimer’s disease (AD) is also growing fast in the current days, and researchers have been trying to understand the mechanisms behind its propagation. Many factors are involved in AD progression, but two protein families, amyloid-beta and tau proteins, are known to be the main contributors [322, 323]. In [45], authors have considered the heterodimer model to incorporate the interaction between these two proteins, each consisting of healthy and toxic densities. Also, they derived the network model corresponding to the PDE model to apply it to AD progression in the brain connectome. Authors in [46] modified the heterodimer model by introducing nonlocal interactions into the model. They studied the following nonlocal network model
$\frac{du_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}u_{k}+u_{j}(a_{0}-a_{1}u_{j})-\frac{a_{2}u_{j}}{1+c_{u}u_{j}}\Phi_{j}\ast\widetilde{u}_{j},$ (43)
$\frac{d\widetilde{u}_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}\widetilde{u}_{k}-\widetilde{a}_{1}\widetilde{u}_{j}+a_{2}\widetilde{u}_{j}\Phi_{j}\ast\bigg{(}\frac{u_{j}}{1+c_{u}u_{j}}\bigg{)},$
$\frac{dv_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}v_{k}+v_{j}(b_{0}-b_{1}v_{j})-b_{3}\widetilde{u}_{j}v_{j}\widetilde{v}_{j}-\frac{b_{2}v_{j}}{1+c_{v}v_{j}}\Phi_{j}\ast\widetilde{v}_{j},$
$\frac{d\widetilde{v}_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}\widetilde{v}_{k}-\widetilde{b}_{1}\widetilde{v}_{j}+b_{3}\widetilde{u}_{j}v_{j}\widetilde{v}_{j}+b_{2}\widetilde{v}_{j}\Phi_{j}\ast\bigg{(}\frac{v_{j}}{1+c_{v}v_{j}}\bigg{)},$
where $u_{j}$ and $\widetilde{u}_{j}$ are the densities of healthy and toxic amyloid beta, and $v_{j}$ and $\widetilde{v}_{j}$ are the densities of healthy and toxic tau proteins at the node $j$ in a given graph $G$. Here, $N$ is the total number of nodes, and $L_{ij}$ are the entries of the Laplacian matrix for the given network [45, 46]. Here, we define the convolution term for a given graph $G$ [46]. For a fixed node $j$ in the network, let $S_{j,1}$ be the set containing all directly connected nodes to the node $j$. We then find the nodes in the graph that are connected to the nodes in the set $S_{j,1}$ except for the node $j$, and we refer to this set as $S_{j,2}$. This process continues until we find the complete set of nodes as $S_{j,1},\ldots,S_{j,m_{j}}$, and $S_{j}$ is the union of this complete set. Following the approach outlined in [46], we find the weight vector $V_{j}$ of $N$ elements, and the convolution term for the node $j$ becomes:
$\Phi_{j}\ast\widetilde{u_{j}}=\sum_{n=1}^{N}V_{j}(n)\widetilde{u}_{n}.$
We can similarly compute the other convolution terms used in the model (43).
There are nonlocal mathematical models on neural networks, especially convolutional neural networks. The traditional neural network blocks feature representations in a local sense; however, long-range dependencies have significant practical learning problems. Here, we describe a nonlocal network in the context of traditional image classification tasks and compare different neural networks. Suppose $X=[X_{1},\ldots,X_{M}]$ is an input sample with $X_{j}(j=1,\dots,M)$ as the feature at position $j$. A nonlocal output signal at position $j$ can be defined as [324]:
$Z_{j}=X_{j}+\frac{W_{Z}}{\mathcal{C}_{j}(X)}\sum_{k}\omega(X_{j},X_{k})g(X_{k}),$
${}Z^{n+1}:=Z^{n}+\mathcal{F}(Z^{n};W^{n}),\leavevmode\nobreak\ \leavevmode\nobreak\ n=0,1,\ldots,N,$ (44)
Genetic regulatory networks (GRNs) are important in coordinating gene expression and orchestrating complex developmental processes in cells. These networks involve transcription factors binding to particular DNA sequences, cis-regulatory elements, and other substances, resulting in a dynamic and linked web of regulatory interactions. Understanding the abnormal alterations in GRNs gives insight into several disease pathways, such as cancer and developmental disorders [325]. Stochastic transcriptional regulation adds a layer of complexity to GRNs, altering cellular heterogeneity, phenotypic variability, and the general adaptability of biological systems [326, 327, 328]. Furthermore, pattern memory in GRNs is important in cellular decision-making processes, allowing cells to respond to the same stimuli differently based on their experience.
The spatio-temporal dynamics of biomolecular networks and biochemical systems are related to the complex dynamic patterns of molecular interactions that occur both in space (inside cellular compartments) and over time [329]. The spatial organization impacts many biomolecular interactions in various cellular compartments, such as the nucleus, cytoplasm, and organelles. Furthermore, spatio-temporal dynamics in metabolic networks are critical for coordinating energy generation, nutrient use, and maintaining cellular homeostasis [330]. Nonetheless, spatial gradients of signalling molecules contribute to generating concentration gradients, impacting cell fate and pattern formation throughout development. The dynamics of transcriptional regulation involve both local and nonlocal components, reflecting the complex interplay of molecular processes within a cell and interactions that extend beyond immediate neighbours. Local dynamics often relate to nearby activities, such as interactions between genes and transcription factors. In contrast, nonlocal dynamics involve events that affect distant genomic areas or operate across distant locations and temporal dimensions.
Nonlocal models are effective for understanding complex dynamics that involve long-range interactions. However, their complexity and computational requirements can sometimes restrict their practical application. It is important to balance accuracy and feasibility when using these models. Here, we provide an overview of the mathematical models discussed so far by highlighting their strengths and limitations:
Models Advantages Limitations
RD model (1) It has a wide range of applications, such as in chemical reactions, biological pattern formation, and even ecological systems The assumption of continuity and homogeneity might not be valid in all mediums such as cellular or molecular discreteness
Fisher-KPP model (3) This model describes the system where interactions are not limited to immediate neighbours but extend over a broader range The range and the mathematical form of the nonlocal interactions can be difficult to measure from empirical observations
Aggregation model (4) The aggregation and depletion of populations can be captured by these models It may not be appropriate for systems where local interactions dominate
Space-time model (5) It describes distant interactions and the delayed responses of populations that shift in space and time High-dimensional integrals increase the computational costs and longer simulation times
Agent-based model (6) It gives insights into complex biological systems through the simulation of individual agents Analyzing the nonlocal model analytically is challenging
Dispersal model (7) It can accurately represent situations where entities are spread out over long distances, such as the spread of diseases It is specific to the particular system and can not be generalized to other environments
Fractional model (9) It adds memory effects, allowing these models to explain systems where the current state depends on the overall history of the system Establishing the suitable fractional order and other parameters in these models can be difficult; especially for the noise or sparse experimental data
Epidemio-logical model (12) These models accurately represent disease distribution across regions and countries, which is critical for understanding global pandemics and epidemics Extensive data on long-distance interactions, travel patterns, and network structure are required to parameterize nonlocal models effectively
COVID-19 model (13) It can capture the spatial distribution of infections while the model (12) does not Obtaining high-quality data on population distribution and mobility patterns
Time-relaxation model (14) It represents the nonlocal dependency of the latent period in disease spread to consider memory effects; as a result, the system’s past states influence its current state It has computational and implementation challenges, especially for large complex systems, and one can avoid them if memory effects are minimal
Age-dependent model (16) These models can take care of age-dependent death and diffusion rates The accuracy of the model is heavily dependent on the quality of the data used and inaccurate data can lead to unreliable model predictions
Cell population model (18) These models can incorporate the effect of adhesion between cells on their movement Deriving conditions for boundedness is substantially harder in higher dimensions than in one dimension
Resource-dependent model (20) These models represent crop raiding effects of large-bodied mammals living in the biodiversity-rich tropics These approaches can not be applied if the mammals have sufficient resources everywhere in the spatial domain
Intra-specific competition model (21) It offers the intra-specific competition of the prey individuals when they compete for resources over larger distances The complexity increases considerably to account for interactions over a broader range of nonlocal interactions, which raises computational demands
Intra-specific-pest competition model (22) These models account for the nonlocal intra-specific competition among prey individuals and pest-type nonlocal interaction with other species The complexity increases considerably to account for interactions over a broader range of nonlocal interactions, which raises computational demands
Shadow model (23) This process reduces the complexity of the original reaction-diffusion model by eliminating the fast-diffusing component and makes the system easier to analyze mathematically This reduction may miss transient behaviours, such as wave propagation or oscillations, that could be important in certain biological contexts
Competition model (24) It can incorporate the nonlocal intra- and inter-specific competitions of the species Nonlocal interaction terms introduce additional parameters related to the range and strength of the nonlocal effects, which can be difficult to estimate from empirical data
Hyperbolic-parabolic model (25) This model accounts for the predators’ flow towards the prey population The choice of the velocity function is crucial and depends on the nature of the predator-prey interactions
Mean field approach (26) This model incorporates the phenomena that the reaction between two chemical particles depends on the reaction radius The effective reaction radius can change with varying temperature, pressure, or medium
Models in game theory (27) and (28) These models can capture the long-range spatial heterogeneity of the strategy space Large strategy space increases the complexity and computational cost of the model
Models in elasticity (29) This theory is highly effective in modelling materials at small scales (e.g., nanomaterials) where long-range atomic interactions significantly affect mechanical behaviour Finding appropriate kernels related to strains and stress is not straightforward and often requires empirical data
Model in materials and structures (34) It can naturally handle discontinuities for modelling materials and structures (e.g., cracks and fractures) without requiring special techniques such as introducing singularities It is challenging when microcrack growth isn’t determined by local deformation or stress
Initial and boundary value problems (35) - (36) These models can incorporate the measurements at more spatial and time points The mathematical and computational parts of these problems are challenging, including the weights related to long-range interactions
Physics-informed neural network (37) It can integrate the physical laws into the training process and estimate the parameter values involved in the laws This approach struggles with problems that have discontinuities, sharp gradients, or complex geometries
Non-equilibrium model (38) This model can be simplified and mathematically tractable and can reduce the computational cost Only low carrier densities and small electric fields allow for the rigorous justification of these drift-diffusion models
Multispecies aggregation model (40) These models incorporate collective behaviours, including swarming, flocking, and pattern creation, which are important in understanding how animals or agents act in groups Reaction terms need to include for capturing quantities change over time due to internal dynamics, such as population growth, competition, predation, etc.
High-fidelity model (41) Optimisation-based approaches enable the model to learn constitutive laws from data by reducing the differences between predicted and observed patterns These methods can suffer from convergence issues, particularly in high-dimensional or highly nonlinear problems
Multiscale and multifidelity model (42) It reduces the overall workload without significantly sacrificing accuracy by using low-cost surrogates for most computations and high-fidelity models when necessary Choosing the appropriate surrogate model is not straightforward, which can impact substantially model accuracy, performance, and computational efficiency
Nonlocal network model (43) It offers a more comprehensive understanding of the structural connectivity and functions of the brain Integrating data from different sources (e.g., MRI, fMRI, PET scans) into nonlocal network models can be difficult
Model in neural network (44) This nonlocal model can quantify the characteristics of the damping effect of nonlocal blocks in a network and is robust to the number of nonlocal blocks Due to the learn long-range features in the network; it needs extra work to study the initialization strategy or to fine-tune the parameters
@@ 3.1 Ecology and epidemiology
The classical Fisher-KPP model and its extensions have been used in many applications, including cell biology, high-energy physics, statistical physics, chemistry, ecology, and epidemiology [331]. Along with the Fisher-KPP model, different types of nonlocal models have been discussed in this article, and we have summarised their applications in this subsection.
@@
In ecological systems, individuals within a population often interact nonlocally due to resource competition, predator-prey interactions, or spatial dispersal. Nonlocal models have been applied to study population dynamics, species coexistence, and spatial patterns in ecological communities. The classical Fisher-KPP equation and its extensions describe the space-time evolution of population densities [80, 81, 332, 333]. The nonlocal Fisher-KPP equation (3) is called the “competitive Lotka-Volterra model” [334, 335, 336]. It is a simple nonlocal model with many applications, e.g., nonlocal intra-specific competition in the predator-prey model [88, 90, 92, 96, 337, 101, 97, 98, 131], global consumption of resources [338], fear effects of prey species [339], cooperative behaviour [340]. As mentioned earlier, the homogeneous solution $u=1$ can be unstable due to nonlocal interaction. For this, we linearize the nonlocal model (3) around the homogeneous solution $u=1$, and then applying the Fourier transform on both sides, we obtain the eigenvalue equation
${}\lambda(k)=-dk^{2}-\widehat{\phi}(k),$ (45)
@@
where $k$ is the wave number and $\widehat{\phi}(k)$ is the Fourier transform of the kernel function $\phi$ [72, 91]. If we choose the kernel function as the Dirac $\delta$-function, then the nonlocal model (3) reduces to the local model (2), and the homogeneous solution $u=1$ becomes stable as $\lambda<0$ for all positive $k$ and $d$. The eigenvalue $\lambda$ may not maintain the negative sign for the other choices of the kernel functions. For example, suppose $\phi$ is a top-hat kernel function [88, 341], then the eigenvalue equation (45) becomes
${}\lambda(k)=-dk^{2}-\frac{\sin(k\delta)}{k\delta},$
@@
and it is not always negative for all $k>0$; rather $\lambda$ can make positive for some $k>0$ by choosing some specific value of $\delta$ and $d$ [341]. For such choices of $\delta$ and $d$, the homogeneous solution $u=1$ becomes unstable, and the nonlocal model produces a non-homogeneous stationary solution. This type of linear stability analysis around the non-trivial homogeneous solution(s) predicts the existence of heterogeneous solution(s) for the nonlocal model, and this method is known as Turing instability analysis [3]. This technique can be applied to the other nonlocal models described in this review article if the homogeneous solution satisfies the Turing instability conditions [4].
@@
The considered nonlocal model with top-hat kernel function has important characteristics. For a fixed value of $\delta$, we may find the critical values $k_{c}$ and $d_{c}$ by which $\lambda=0$ holds for unique $k=k_{c}$ and $d=d_{c}$, and $\lambda>0$ holds for a range of values of $k$ with $d<d_{c}$. However, these critical values do not exist for small values of $\delta$ as $\lambda$ negative for $\delta$ approaches 00. Hence, $\delta$ (range of nonlocal interaction) requires a minimum value to unstable the homogeneous solution $u=1$ [88, 92]. In addition, two length scales can be identified in the presence of the top-hat kernel function: (i) the diffusion parameter $d$ as activation and (ii) the range of nonlocal interaction $\delta$ as inhibition. We can observe a Turing instability for $d\ll\delta$ (short-range activation and long-range inhibition), and it produces a Turing pattern [3, 35, 342]. On the other hand, for $\delta\ll d$ (long-range activation and short-range inhibition), travelling wavefront propagation can be observed. In particular, for $\delta\rightarrow 0$, the nonlocal model reduces to the local classical Fisher-KPP equation and exhibits a travelling wave solution [4, 343]. The same type of behaviours can be observed for other truncated kernel functions (triangular, parabolic, etc.), but the minimum range of the nonlocal interaction varies for different functions [92].
@@
In plant communities, the local competition term could represent the competition for light, assuming the canopy size is negligible compared to the range of the root system [209]. If the opposite holds, nonlocal competition comes into the picture [212, 344]. In this case, individuals compete to access the resources, which can redistribute itself [47, 345]. In ecology, nonlocal models can capture many factors; e.g., Cantrell and Cosner [346] constructed a model for ladybirds feeding on aphids. They focused on the nonlocal RD model for aphid density in a particular patch that involves the whole population.
@@
Another type of solution called ‘travelling wave solution’ is studied for a partial differential equation that propagates with a constant speed while maintaining its shape in space. These solutions represent wave-like patterns that propagate through space while maintaining their shape and speed. Travelling wave solutions are common in biology, and numerous models in population dynamics produce biological waves. In the context of population dynamics, the travelling wave manifests itself as a wave of change in population density through a habitat, e.g., a plague that travels through a continent. This travelling wave solution also signifies the invasion of the species over the habitat. The travelling wave solutions have been studied for the nonlocal models [347, 348], e.g., let us consider an extension of the nonlocal FKPP model (3) as:
@@
The initial condition plays a crucial role in finding the travelling wave solutions. A small population is introduced in a localized space, which is considered the initial condition for finding the solutions to the travelling wave. In addition, the boundary conditions for this nonlocal model are considered as $u_{x}(0,t)=0$ and $u\rightarrow 0$ for $x\rightarrow\infty$ [347]. For the “thin-tailed” kernel function $\psi(z)=\exp{(-|z|)}/2$, the nonlocal model (46) can be converted to the coupled equations [347]:
@@
with localized initial conditions for the species $u$ and boundary conditions $u_{x}(0,t)=0$, $u\rightarrow 0$, $w_{x}(0,t)=0$, and $w\rightarrow 0$ for $x\rightarrow\infty$ [347]. Sometimes, this conversion technique can be applied to the nonlocal model with space-time nonlocality [92], e.g., the single species model considered in (5). Along with travelling wave solutions, periodic travelling waves (or wavetrain) are also observed in local and nonlocal models, which are the periodic solutions of travelling waves. These travelling wave solutions also signify the migration of two or more sub-populations.
The single species nonlocal model (4) in population dynamics was introduced by Britton [99, 100], and it is an advanced version of the local FKPP model as it can be obtained by considering $\phi$ as the delta function. The nonlocal model (4) includes local aggregation that takes care of animal grouping for protective measures against predation, which could arise for many reasons, such as grassland herds, schools of fish or flocks of birds. Sometimes, species take time to move from one position to another. In this case, the nonlocal model (5) containing a spatio-temporal average weighted toward the current time and position better fit these types of scenarios [100]. This nonlocal model accounts for the influence of past conditions on present ecological processes.
Nonlocal interactions are important in multi-body systems because each body’s behaviour is influenced by its nearby neighbours and bodies at a distance. Such interactions are common in a wide range of physical and biological systems, including granular materials, colloidal suspensions, and biomolecular systems. In the case of population biology, species sense their surroundings to make different decisions, called nonlocal sensing, and these happen on different scales [48, 49, 293]. The nonlocal sensing effect in population spatial distributions is modelled by the equations (39) and (40) and leads to species aggregation, segregation, or complex patterns [294, 295, 296, 297]. In addition, the nonlocal multi-species advection-diffusion model (40) can capture a wide variety of patterns [300, 305]. These types of models have been applied to predator-prey interactions [349], cell-sorting [350], and animal territoriality [351]. Authors in [300] showed that the inclusion of self-attractive terms causes oscillatory solutions for the two-species model, and a qualitative effect on patterns occurs due to a change in the width of the spatial average.
@@
The nonlocal dispersal used in the model (7) is important for plants and animals. It characterizes the movements of organisms in a habitat. For plants, this nonlocal dispersal happens through the seed dispersal process. In the review article, Bullock et al. fitted 11 types of probability density functions as dispersal kernel functions for 144 plant species, and all of them are good fits to the grouped data sets [352]. In addition, the dispersal-competition models have been used in the integrodifference equations, which are discrete in time and continuous in space [353]. For instance, the integrodifference equations of a population size $n(t,x)$ at time $t$ and position $x$ can have the form [353]:
@@
where $\phi(x,y)$ is the kernel function that satisfies the normality condition over the spatial domain $R$. In the absence of dispersal, the growth equation is of the form $n(t+1,x)=f(n(t,x))$. These models apply to the development and spread of seed-dispersed plants. Moreover, a similar model can be considered for the single-species annual plant model when seeds do not survive for more than one year [353]. In addition, the nonlocal dispersal is also applied to study the metapopulation dynamics. In [354], the author studied the effects of nonlocal dispersal on population persistence and the conservation of species in fragmented habitats where local populations are interconnected through dispersal, colonization, and extinction processes. It has implications for ecosystem resilience, invasive species management, and conservation biology. As research in ecology and population dynamics progresses, the study of nonlocal dispersal will continue to shed light on the intricate processes that shape biodiversity and ecological patterns.
@@
In [355], Hastings et al. provide insights into the spatial spread of invasions, a specific example of nonlocal dispersal in ecology. The paper discusses theoretical developments and empirical evidence related to the spatial dynamics of invasive species, highlighting the significance of nonlocal dispersal in driving the spread of invasive populations. Furthermore, in [356], the authors delve into the role of diffusion in ecological problems. It covers various aspects of dispersal and diffusion in ecology, including nonlocal dispersal and its effects on population dynamics and spatial ecology. The book provides a comprehensive and interdisciplinary perspective on diffusion in ecological systems and is an excellent resource for those interested in applying nonlocal dispersal in biology.
@@
These applications demonstrate the versatility of fractional nonlocal models in capturing the complexities of biological systems. Fractional calculus provides a powerful framework to describe anomalous transport, long-range interactions, and memory effects, offering valuable insights into various biological processes. For instance, FODEs are used to study the memory effect on population growth or some inherent randomness at a microscopic level that regulates the macroscopic dynamics [357]. In ecology, FODEs have been studied in predator-prey models to account for the interactions of the population densities dependent on a certain time in the past [358, 359, 360]. The nonlocal FODE model (9) describes the emergence and evolution of biological species in the presence of memory. Furthermore, the fractional integral representation is used in the statistical thermodynamics of confined systems to obtain generalized transport equations, distribution functions, and correlation functions that take into account the particular characteristics of confinement-induced dynamics [361].